# Zero shot Generalization of Vision-Based RL Without Data Augmentation

### v1

@@ 1 Introduction
Training generalist agents that can adapt to novel environments and unseen task variations is a longstanding goal for vision-based RL. RL generalization benchmarks have focused on data augmentation to increase the amount of training data available to the agent while preventing model overfitting and increasing robustness to environment perturbations (Yarats et al., 2021a; Almuzairee et al., 2024; Hansen et al., 2021). This follows the current trend in the broader robot learning community of training large models at scale on massive datasets (Kim et al., 2024; Hansen et al., 2024; Team et al., 2024) with the hope that the model will generalize. However, a significant drawback of these approaches is, intuitively, that they require larger models, more training data, longer training times, and have greater training instability that must be dealt with care when training RL agents.
Yet when we examine biological agents, we find that humans and, indeed, many other primates are able to quickly adapt to task variations and environment perturbations DiCarlo et al. (2012); Friston (2010). While all aspects of biological intelligence that contribute to generalization have yet to be understood, there is some understanding in the recent neuroscience literature of aspects related to representation learning that we look to for inspiration. Many parts of the brain in human and non-human primates contain neurons that represent single factors of variation within the environment, such as grid cells (Hafting et al., 2005), object-vector cells (Høydal et al., 2019), and border cells (Solstad et al., 2008) that represent euclidean spaces, distance to objects, and distance to borders, respectively. Such disentangled representations have been theorized to facilitate compositional generalization (Higgins et al., 2018) and have been studied with curated datasets where the factors of variation are known (Higgins et al., 2017a; Whittington et al., 2023), and even within the context of RL (Higgins et al., 2017b). It is then to our surprise, with limited exceptions (Dunion et al., 2023; Sax et al., 2018), that disentangled representation learning has not garnered much attention within robot learning or RL more generally. One potential reason for this is that learning disentangled representations while simultaneously learning an RL policy is extremely difficult. Indeed, Higgins et al. (2017b) required a two-stage approach where the disentangled representation was learned first, followed by policy learning. In addition, Yarats et al. (2021b) found that using a $\beta$-VAE directly led to training instability and worse performance, instead opting to use a deterministic autoencoder with softer constraints. Finally, there is counter-evidence by Schott et al. (2022) to suggest that, while disentangling representations may facilitate generalization, it alone cannot achieve out-of-distribution (OOD) generalization.
@@
We hypothesize that one of the potential key missing ingredients to OOD generalization is associative memory mechanisms that use prior experiences to help inform decision-making in light of new data in a disentangled latent space. Intuitively, if the representation of a stored memory and new observation are disentangled, then the projection of the new observation onto the stored memories becomes a factorized projection in which individual factors can be compared independently of other factors (Figure 1). Indeed, many of the aforementioned single-factor neurons exist in the entorhinal cortex in the hippocampus (Hafting et al., 2005), responsible for, amongst other things, memory recollection and association. Recent literature suggests that the hippocampus learns flexible representations of memories by decomposing sensory information into reusable components and has been implicated in other cognitive tasks such as planning, decision-making, and imagining novel scenarios (Behrens et al., 2018; Rubin et al., 2014). It would then seem that the disentanglement of high-dimensional data into a modular, reusable representation is simply the first step in a multi-step process that enables generalization in biological agents. Inspired by these two ingredients found in nature, we propose a new method, ALDA (Associative Latent DisentAnglement), that 1. learns a disentangled representation from the training data and 2. uses an associative memory model to recover data points in the original training distribution zero-shot given OOD data. We demonstrate how this approach enables zero-shot generalization on a common generalization benchmark for vision-based RL without using data augmentation techniques or techniques that remove distractor variables from the latent space. Finally, we provide a formal proof showing that data augmentation methods for vision-based RL create what we refer to as "weak" disentanglement, where the latent space is perhaps partitioned into two or more categories but not perfectly factorized into individual subcomponents. We conclude by discussing the implications of this insight and future directions of this line of research.
@@ 2.1 Reinforcement Learning
We wish to learn a policy $\pi$ that maps states to optimal actions that maximize cumulative reward. The agent-environment interaction loop is typically formulated as a Markov Decision Process (MDP) $(\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{P},\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ are the state and action spaces, $\mathcal{R}(s,a)$ is the reward function, $\mathcal{P}(s^{\prime}|s,a)$ is the probabilistic transition function, and $\gamma$ is the discount factor. The policy $\pi$ learns a mapping of state to action with the objective of maximizing cumulative discounted return $G_{t}=\mathbb{E}\left[\sum_{t=0}^{T}\gamma^{t}R(s_{t},a_{t})\right]$. In vision-based RL, we do not assume access to the low dimensional state $s_{t}\in\mathcal{S}$. Instead, we must infer $s_{t}$ given high-dimensional image observations $o_{t}\in\mathcal{O}$, making the problem a partially observable MDP, or POMDP $(\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{P},\mathcal{O},\gamma)$, where $\mathcal{O}$ is the space of high-dimensional observations.
@@
$J(Q)=\mathbb{E}_{(s_{t},a_{t},s_{t+1}\sim\mathcal{D}}\left[(Q(s_{t},a_{t})-r_{t}-\gamma y)^{2}\right].$
Here, $y$ is the soft $Q$-target, which is computed as $y=r(s_{t},a_{t})+\gamma(\text{min}_{\theta_{1,2}}Q_{\theta_{i}^{{}^{\prime}}}(s_{t+1},a_{t+1})-\alpha\text{log}\pi(\cdot|s_{t+1}))$. We can then describe the policy’s objective as:
$J(\pi)=\mathbb{E}_{s_{t}\sim\mathcal{D}}\left[\text{min}_{i=1,2}Q_{\theta_{i}}(s_{t},a_{t})-\alpha\text{log}\pi(a_{t}|s_{t})\right].$
@@
where $g:\mathcal{S}\rightarrow\mathcal{O}$ is the non-linear data generating function. The nonlinear ICA problem is to recover the underlying sources given samples from this model. Similarly, the goal of latent disentanglement is to learn a latent representation z such that every variable $_{1},...,z_{n_{s}}\in\textbf{z}$ z corresponds to a distinct source $s_{1},...,s_{n_{s}}$. Unfortunately, nonlinear ICA is nonidentifiable – that is, there are many decompositions of the data into sets of independent latents that fit the dataset, and so recovering the true sources reliably is impossible. Thus, the field of disentangled representation learning has focused more on empirical results and evaluation metrics on toy datasets where the true sources of variation are known. Given a dataset of paired source-data samples ${(\textbf{s},\textbf{o}=g(\textbf{s}))}$, the goal is to learn an encoder $f:\mathcal{O}\rightarrow\mathcal{Z}$ and a decoder $\hat{g}:\mathcal{Z}\rightarrow\mathcal{O}$ such that the disentanglement evaluation metrics are high while also maintaining acceptable reconstructions of the data. Disentanglement models are typically constructed as (variational) autoencoders (Whittington et al., 2023; Hsu et al., 2023; Higgins et al., 2017a) and are rarely applied outside of toy datasets.
@@ 2.3 Generalization in Vision-Based RL
Image augmentation methods have shown success and have become the go-to method for generalizing vision-based RL algorithms such as Soft Actor-Critic (SAC) (Haarnoja et al., 2018) and TD3 (Fujimoto et al., 2018), generally using augmentations such as random crops, random distortions, and random image overlays to simulate distracting backgrounds. Methods such as DrQ (Yarats et al., 2021a), SADA (Almuzairee et al., 2024), and SVEA (Hansen et al., 2021) regularize the $Q$ function by providing the original and augmented images as inputs into the critic. In many cases, however, the image augmentations can put the training data within the support of the distributions of the evaluation environments. For example, the "random convolution" image augmentation changes the color of the agent and/or background, and the policy is evaluated on an environment where the color of the agent is randomized. This brings into question whether these methods are truly capable of zero-shot extrapolative generalization when the training data is made to be sufficiently similar to the test data.
@@
Beyond image augmentation techniques are methods that perform self-supervision using auxiliary objectives. Note that image augmentations for RL are also sometimes referred to as self-supervised objectives, however we wish to make the distinction between methods that leverage data augmentation and those that don’t. DARLA Higgins et al. (2017b), to the best of our knowledge, is the only prior method that learns a disentangled representation of the image inputs using a highly regularized $\beta$-VAE (Higgins et al., 2017a) for zero-shot generalization in RL. DARLA’s approach is two-stage, where an initial dataset is collected by sampling random actions to first learn a disentangled latent representation, and then a policy is trained on this representation to maximize future return. However, a significant shortcoming is that random actions may not cover the full state distribution of the agent for more complicated tasks, whereas our method jointly learns the disentangled representation and the policy. SAC+AE (Yarats et al., 2021b) trains a decoder to reconstruct the images, resulting in a rich latent space that improves performance and sample efficiency on many vision-based tasks. Interestingly, SAC+AE mentions $\beta$-VAE’s used by DARLA and proposes using a deterministic variant with similar constraints that shows some zero-shot generalization capability, but the authors make no mention of disentanglement and instead conclude that the key ingredient was adding a reconstruction loss as an auxiliary objective.
@@
Another promising approach to generalization is learning a task-centric or object-centric representation using auxiliary objectives. Yamada et al. (2022) learn a task-centric representation by using expected discounted returns as labels, with the auxiliary task being to minimize the error between the predicted and true return values using the learned representation. Ferraro et al. (2023); Pore et al. (2024) use segmentation masks to learn object-centric representations that are robust to background distractors. One drawback of these approaches is that the latent representation overfits to the task by excluding all other information not relevant to the immediate task, usually citing that irrelevant information in the latent space hinders generalization performance. However, adapting to a new task that involves information that was previously considered irrelevant becomes a challenge for these methods. We hypothesize that the issue is not having "irrelevant" information in the latent space but rather that the latent variables are entangled without strong priors for disentanglement. A disentangled representation then paves the way for association, whereby individual dimensions of latent vectors from OOD images can be independently zero-shot mapped back to known values of those latent variables learned from the training data.
@@ 3 On the Relationship Between Disentanglement and Data Augmentation
We begin by motivating the case for learning a disentangled representation for RL agents by showing a connection between data augmentation and disentangled representation learning. Specifically, we formally prove that data augmentation is a weak disentanglement of the latent space. We define weak disentanglement as a partial factorization of some but perhaps not all latent dimensions of the latent space. Strong disentanglement, on the other hand, is a complete factorization where each latent dimension $_{k}\in\textbf{z}$ z encodes for a unique source $_{i}\in\textbf{s}$ s and is thus linearly independent of other latent dimensions, which is the goal of disentangled representation learning. The full proof is provided in A.1.
Theorem 1: Suppose we are given a $\textbf{z}=f_{\theta}(g(\textbf{s}))$, where some latent dimension $_{k}\in\textbf{z}$ z approximates one or more sources. We will denote the approximations as $\hat{s_{i}}$. We can categorize the sources s into two categories, $D$ and $E$, which correspond to task-relevant and task-irrelevant sources, respectively. For any such $z_{k}$, if $Q^{*}(\textbf{z},a)$ is an optimality invariant optimal Q-function immune to distractor variables, then the following must be true of z:
$cov(\hat{s_{i}},\hat{s_{j}}|z_{k})=0\text{ }\forall s_{i}\in D,s_{j}\in E,z_{k}\in\textbf{z}.$ (2)
@@
$p(s_{1},...,s_{k})=\sum_{s_{k+1}}\times\sum_{s_{k+2}}\times...\times\sum_{s_{n_{s}}}p(s_{1},...,s_{k},s_{k+1},...s_{n_{s}}).$ (3)
The implication of 3 is that we must collect data for every possible variation of the task-irrelevant sources, which may be prohibitively expensive for real-world applications. Instead, a model with strong priors that leads to disentanglement without data augmentation essentially achieves the same result without the additional costs. Although the latent representation of such a model may still contain task-irrelevant features, the policy can learn to simply ignore them or associate them with known values in the presence of OOD data, as is the case with ALDA. In addition, these task-irrelevant features may become relevant if the task changes (e.g., if the current task is for a manipulator to stack a blue cube, and the next task is to stack a red cube), and so it may, in fact, be important to keep them.
@@ 4.1 Disentanglement
We now describe our framework for jointly learning a disentangled representation and performing association. For latent disentanglement, we choose to use QLAE Hsu et al. (2023), the current SOTA disentanglement method, which trains an encoder $f_{\theta}$ that maps to a continuous disentangled latent space, a discrete, parameterized latent model $l_{\psi}$, and a decoder $g_{\phi}$ that reconstructs the observation. Similar to VQ-VAE (van den Oord et al., 2017), QLAE uses a discrete codebook for the latent space, except that each dimension uses its own separate scalar codebook. Concretely, $Z$ is the set of latent codes defined by the Cartesian product of $n_{z}$ scalar codebooks $Z=V_{1}\times...\times V_{n_{z}}$ i.e., there are $n_{z}$ latent variables and $|V_{j}|$ discrete categories per variable. The continuous outputs of the encoder are the latent variables, each of which is quantized to the nearest scalar value in their respective codebooks.
@@
Since we cannot differentiate through $argmin$, as with VQ-VAE, the authors of QLAE use quantization and commitment losses and a straight-through gradient estimator (Bengio et al., 2013):
$\mathcal{L}_{\text{quantize}}=||\text{StopGradient}f_{\theta}(x))-z_{d}||_{2}^{2},\text{ }\mathcal{L}_{\text{commit}}=||f_{\theta}(x)-\text{StopGradient}(z_{d})||_{2}^{2}.$ (5)
The authors claim that while this is a failure mode for vector quantization, $Z$ is low-dimensional enough that, in practice, it does not meaningfully impact performance. While this may be true for standalone disentanglement benchmarks, we find that it causes training instability and performance degradation when jointly learning a policy for high-dimensional continuous control problems We propose a solution in section 4.2 from the viewpoint of associative memory.
It is common practice in many vision-based RL algorithms to utilize framestacking to incorporate temporal information into the latent space. This means that the encoder accepts as input, and the decoder produces a stack of RGB images in $\mathbb{R}^{B\times Ck\times H\times W}$, where $k$ is the number of frames. However, latent disentanglement models have only been shown to work on datasets of singular images and struggle to disentangle sources of individual images when given stacks of images as inputs. Evidence of this is presented in the appendix, Section A.5. To resolve this issue, we fold $k$ into the batch dimension and encode/decode batches of single images in $\mathbb{R}^{Bk\times C\times H\times W}$, resulting in a batch size $Bk$ of disentangled latent vectors $z_{d}\in\mathbb{R}^{Bk\times n_{s_{i}}}$. To incorporate temporal information, we reshape the batch of latent vectors into $\mathbb{R}^{B\times kn_{s_{i}}}$ and feed it into a 1D convolutional neural network (CNN), producing our final latent vector $z\in\mathbb{R}^{B\times e}$. $z$ is used as the state representation for the actor and critic networks, while the decoder network for the disentanglement model only ever receives the disentangled representation $z_{d}$ as input.

### v2

@@ 1 Introduction
Training generalist agents that can adapt to novel environments and unseen task variations is a longstanding goal for vision-based RL. RL generalization benchmarks have focused on data augmentation to increase the amount of training data available to the agent while preventing model overfitting and increasing robustness to environment perturbations (Yarats et al., 2021a; Almuzairee et al., 2024; Hansen et al., 2021). This follows the current trend in the broader robot learning community of training large models at scale on massive datasets (Kim et al., 2024; Hansen et al., 2024; Team et al., 2024) with the hope that the model will generalize.
@@
However, a significant drawback of these approaches is, intuitively, that they require larger models, more training data, longer training times, and have greater training instability that must be dealt with care when training RL agents.
Yet when we examine biological agents, we find that humans and, indeed, many other primates are able to quickly adapt to task variations and environment perturbations (DiCarlo et al., 2012; Friston, 2010). While all aspects of biological intelligence that contribute to generalization have yet to be understood, there is some understanding in the recent neuroscience literature of aspects related to representation learning that we look to for inspiration. Many parts of the brain in human and non-human primates contain neurons that represent single factors of variation within the environment, such as grid cells (Hafting et al., 2005), object-vector cells (Høydal et al., 2019), and border cells (Solstad et al., 2008) that represent euclidean spaces, distance to objects, and distance to borders, respectively. Such disentangled representations have been theorized to facilitate compositional generalization (Higgins et al., 2018) and have been studied with curated datasets where the factors of variation are known (Higgins et al., 2017a; Whittington et al., 2023), and even within the context of RL (Higgins et al., 2017b). It is then to our surprise, with limited exceptions (Dunion et al., 2023; Sax et al., 2018), that disentangled representation learning has not garnered much attention within robot learning or RL more generally. One potential reason for this is that learning disentangled representations while simultaneously learning an RL policy is extremely difficult. Indeed, Higgins et al. (2017b) required a two-stage approach where the disentangled representation was learned first, followed by policy learning. In addition, Yarats et al. (2021b) found that using a $\beta$-VAE directly led to training instability and worse performance, instead opting to use a deterministic autoencoder with softer constraints. Finally, there is counter-evidence by Schott et al. (2022) to suggest that, while disentangling representations may facilitate generalization, it alone cannot achieve out-of-distribution (OOD) generalization.
@@
We hypothesize that one of the potential key missing ingredients to OOD generalization is associative memory mechanisms that use prior experiences to help inform decision-making in light of new data in a disentangled latent space. Intuitively, if the representation of a stored memory and new observation are disentangled, then the projection of the new observation onto the stored memories becomes a factorized projection in which individual factors can be compared independently of other factors (Figure 1). Indeed, many of the aforementioned single-factor neurons exist in the entorhinal cortex in the hippocampus (Hafting et al., 2005), responsible for, amongst other things, memory recollection and association. Recent literature suggests that the hippocampus learns flexible representations of memories by decomposing sensory information into reusable components and has been implicated in other cognitive tasks such as planning, decision-making, and imagining novel scenarios (Behrens et al., 2018; Rubin et al., 2014). It would then seem that the disentanglement of high-dimensional data into a modular, reusable representation is simply the first step in a multi-step process that enables generalization in biological agents.
Inspired by these two ingredients found in nature, we propose a new method, ALDA (Associative Latent DisentAnglement), that 1. learns a disentangled representation from the training data and 2. uses an associative memory model to recover data points in the original training distribution zero-shot given OOD data. We demonstrate how this approach enables zero-shot generalization on a common generalization benchmark for vision-based RL without using data augmentation techniques or techniques that remove distractor variables from the latent space. Finally, we provide a formal proof showing that data augmentation methods for vision-based RL create what we refer to as weak disentanglement, where the latent space is perhaps partitioned into two or more categories but not perfectly factorized into individual subcomponents. We conclude by discussing the implications of this insight and future directions of this line of research.
@@ 2.1 Reinforcement Learning
We wish to learn a policy $\pi$ that maps states to optimal actions that maximize cumulative reward. The agent-environment interaction loop is typically formulated as a Markov Decision Process (MDP) $(\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{P},\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ are the state and action spaces, $\mathcal{R}(s,a)$ is the reward function, $\mathcal{P}(s^{\prime}|s,a)$ is the probabilistic transition function, and $\gamma$ is the discount factor. The policy $\pi$ learns a mapping of state to action with the objective of maximizing cumulative discounted return $G_{t}=\mathbb{E}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}\right]$. In vision-based RL, we do not assume access to the low dimensional state $s_{t}\in\mathcal{S}$. Instead, we must infer $s_{t}$ given high-dimensional image observations $o_{t}\in\mathcal{O}$, making the problem a partially observable MDP, or POMDP $(\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{P},\mathcal{O},\gamma)$, where $\mathcal{O}$ is the space of high-dimensional observations.
@@
$J(Q)=\mathbb{E}_{(s_{t},a_{t},s_{t+1}\sim\mathcal{D}}\left[(Q(s_{t},a_{t})-\gamma y)^{2}\right].$
Here, $y$ is the soft $Q$-target, which is computed as $y=r_{t}+\gamma(\text{min}_{\theta_{1,2}}Q_{\theta_{i}^{{}^{\prime}}}(s_{t+1},a_{t+1})-\alpha\text{log}\pi(\cdot|s_{t+1}))$. We can then describe the policy’s objective as:
$J(\pi)=\text{max }_{\theta}\mathbb{E}_{s_{t}\sim\mathcal{D}}\left[\text{min}_{i=1,2}Q_{\theta_{i}}(s_{t},a_{t})-\alpha\text{log}\pi(a_{t}|s_{t})\right].$
@@
where $g:\mathcal{S}\rightarrow\mathcal{O}$ is the non-linear data generating function. The nonlinear ICA problem is to recover the underlying sources given samples from this model. Similarly, the goal of disentanglement is to learn a latent representation z such that every variable $z_{1},...,z_{n_{s}}\in\textbf{z}$ corresponds to a distinct source $s_{1},...,s_{n_{s}}$. Unfortunately, nonlinear ICA is nonidentifiable – that is, there are many decompositions of the data into sets of independent latents that fit the dataset, and so recovering the true sources reliably is impossible. Thus, the field of disentangled representation learning has focused more on empirical results and evaluation metrics on toy datasets where the true sources of variation are known. Given a dataset of paired source-data samples $(\textbf{s},\textbf{o}=g(\textbf{s}))$}, the goal is to learn an encoder $f:\mathcal{O}\rightarrow\mathcal{Z}$ and a decoder $\hat{g}:\mathcal{Z}\rightarrow\mathcal{O}$ such that the disentanglement evaluation metrics are high while also maintaining acceptable reconstructions of the data. Disentanglement models are typically constructed as (variational) autoencoders (Whittington et al., 2023; Hsu et al., 2023; Higgins et al., 2017a) and are rarely applied outside of toy datasets.
@@ 2.3 Generalization in Vision-Based RL
Image augmentation methods have shown success and have become the go-to method for generalizing vision-based RL algorithms such as Soft Actor-Critic (SAC) (Haarnoja et al., 2018) and TD3 (Fujimoto et al., 2018), generally using augmentations such as random crops, random distortions, and random image overlays to simulate distracting backgrounds. Methods such as DrQ (Yarats et al., 2021a), SADA (Almuzairee et al., 2024), and SVEA (Hansen et al., 2021) regularize the $Q$ function by providing the original and augmented images as inputs into the critic. In many cases, however, the image augmentations can put the training data within the support of the distributions of the evaluation environments. For example, the random conv image augmentation changes the color of the agent and/or background, and the policy is evaluated on an environment where the color of the agent is randomized. This brings into question whether these methods are truly capable of zero-shot extrapolative generalization when the training data is made to be sufficiently similar to the test data.
@@
Beyond image augmentation techniques are methods that perform self-supervision using auxiliary objectives. Note that image augmentations for RL are also sometimes referred to as self-supervised objectives, however we wish to make the distinction between methods that leverage data augmentation and those that do not. DARLA Higgins et al. (2017b), to the best of our knowledge, is the only prior method that learns a disentangled representation of the image inputs using a highly regularized $\beta$-VAE (Higgins et al., 2017a) for zero-shot generalization in RL. DARLA’s approach is two-stage, where an initial dataset is collected by sampling random actions to first learn a disentangled latent representation, and then a policy is trained on this representation to maximize future return. However, a significant shortcoming is that random actions may not cover the full state distribution of the agent for more complicated tasks, whereas our method jointly learns the disentangled representation and the policy. SAC+AE (Yarats et al., 2021b) trains a decoder to reconstruct the images, resulting in a rich latent space that improves performance and sample efficiency on many vision-based tasks. Interestingly, SAC+AE mentions $\beta$-VAE’s used by DARLA and proposes using a deterministic variant with similar constraints that shows some zero-shot generalization capability, but the authors make no mention of disentanglement and instead conclude that the key ingredient was adding a reconstruction loss as an auxiliary objective.
@@
Another promising approach to generalization is learning a task-centric or object-centric representation using auxiliary objectives. Yamada et al. (2022) learn a task-centric representation by using expected discounted returns as labels, with the auxiliary task being to minimize the error between the predicted and true return values using the learned representation. Ferraro et al. (2023); Pore et al. (2024) use segmentation masks to learn object-centric representations that are robust to background distractors. Zhang et al. (2021) utilizes bisimulation metrics to determine behavioral similarity between states, resulting in robust, invariant representations. One drawback of these approaches is that the latent representation overfits to the task by excluding all other information not relevant to the immediate task, usually citing that irrelevant information in the latent space hinders generalization performance. However, adapting to a new task that involves information that was previously considered irrelevant becomes a challenge for these methods. We hypothesize that the issue is not having irrelevant information in the latent space but rather that the latent variables are entangled without strong priors for disentanglement. A disentangled representation then paves the way for association, whereby individual dimensions of latent vectors from OOD images can be independently zero-shot mapped back to known values of those latent variables learned from the training data.
@@ 3 On the Relationship Between Disentanglement and Data Augmentation
[…]
Theorem 1: Suppose we are given a $\textbf{z}=f_{\theta}(g(\textbf{s}))$, where some latent dimension $z_{k}\in\textbf{z}$ approximates one or more sources. We will denote the approximations as $\hat{s_{i}}$. We can categorize the sources s into two categories, $D$ and $E$, which correspond to task-relevant and task-irrelevant sources, respectively. For any such $z_{k}$, if $Q^{*}(\textbf{z},a)$ is an optimality invariant optimal Q-function immune to distractor variables, then the following must be true of z:
[…]
@@
$p(s_{1},...,s_{k})=\sum_{s_{k+1}}\sum_{s_{k+2}}...\sum_{s_{n_{s}}}p(s_{1},...,s_{k},s_{k+1},...s_{n_{s}}).$
The implication of this equation is that we must collect data for every possible variation of the task-irrelevant sources, which may be prohibitively expensive for real-world applications. Instead, a model with strong priors that leads to disentanglement without data augmentation essentially achieves the same result without the additional costs. Although the latent representation of such a model may still contain task-irrelevant features, the policy can learn to simply ignore them or associate them with known values in the presence of OOD data, as is the case with ALDA. In addition, these task-irrelevant features may become relevant if the task changes (e.g., if the current task is for a manipulator to stack a blue cube, and the next task is to stack a red cube), and so it may, in fact, be important to keep them.
@@ 4.1 Disentanglement
We now describe our framework for jointly learning a disentangled representation and performing association. For latent disentanglement, we choose to use QLAE (Hsu et al., 2023), the current SOTA disentanglement method, which trains an encoder $f_{\theta}$ that maps to a continuous disentangled latent space, a discrete, parameterized latent model $l_{\psi}$, and a decoder $g_{\phi}$ that reconstructs the observation. Similar to VQ-VAE (van den Oord et al., 2017), QLAE learns a discrete latent representation, except that each dimension of the continuous representation is mapped to a discrete scalar via a separate codebook, rather than having one codebook for the entire latent space. Concretely, $Z$ is the set of latent codes defined by the Cartesian product of $n_{z}$ scalar codebooks $Z=V_{1}\times...\times V_{n_{z}}$ i.e., there are $n_{z}$ latent variables and $|V_{j}|$ discrete categories per variable. The continuous outputs of the encoder are the latent variables, each of which is quantized to the nearest scalar value in their respective codebooks i.e. the number of codebooks equals the dimensionality of the latent representation.
@@
Since we cannot differentiate through argminargmin, as with VQ-VAE, the authors of QLAE use quantization and commitment losses and a straight-through gradient estimator (Bengio et al., 2013):
$\mathcal{L}_{\text{quantize}}$ $=||\text{StopGradient}f_{\theta}(x))-z_{d}||_{2}^{2},$ (4)
[…]
The authors claim that while this is a failure mode for vector quantization, $Z$ is low-dimensional enough that, in practice, it does not meaningfully impact performance. While this may be true for standalone disentanglement benchmarks, we find that it causes training instability and performance degradation when jointly learning a policy for high-dimensional continuous control problems. We propose a solution in section 4.2 from the viewpoint of associative memory.
In vision-based RL, framestacking is commonly used to encode temporal information by stacking $k$ RGB frames, resulting in encoder inputs and decoder outputs of shape $\mathbb{R}^{B\times Ck\times H\times W}$. However, latent disentanglement models perform poorly when given stacked images, struggling to separate independent sources. We observe this in Appendix Section A.5. To resolve this issue, we fold $k$ into the batch dimension and encode/decode batches of single images in $\mathbb{R}^{Bk\times C\times H\times W}$, resulting in a batch size $Bk$ of disentangled latent vectors $z_{d}\in\mathbb{R}^{Bk\times n_{s_{i}}}$. To incorporate temporal information, we reshape the batch of latent vectors into $\mathbb{R}^{B\times kn_{s_{i}}}$ and feed it into a 1D convolutional neural network (CNN), producing our final latent vector $\in\mathbb{R}^{B\times e}. zz$ is used as the state representation for the actor and critic networks, while the decoder network for the disentanglement model only ever receives the disentangled representation $z_{d}$ as input.